OpenStudy (narissa):

what is the slope of the linear function table shows x with a line next to it ,-2,1,2,4 y 13,1,-3,-11

OpenStudy (narissa):

@iGreen

OpenStudy (narissa):

@Rushwr

OpenStudy (narissa):

is it 1?

OpenStudy (narissa):

@ChantySquirrel1129**

OpenStudy (igreen):

Picture?

OpenStudy (igreen):

Take any 2 points and plug them into the slope formula.

OpenStudy (igreen):

(1, 1), (2, -3) x1 y1 x2 y2 \(\sf m = \dfrac{y_2-y_1}{x_2-x_1}\) \(\sf m = \dfrac{-3-1}{2+1}\)

OpenStudy (narissa):

2/3

OpenStudy (narissa):

-2/3

OpenStudy (igreen):

My bad, that should be 2 - 1

OpenStudy (narissa):

-2/1

OpenStudy (igreen):

Check again..-3 - 1 isn't -2

OpenStudy (narissa):

-4

OpenStudy (narissa):

-4/1

OpenStudy (igreen):

Correct. So -4 is our slope